# Zining Cheng and Guanzhou Ji

### v1

@@ Abstract
Keywords: HDR Photography, 3D Estimation, Panorama, Lighting Simulation
@@ 1 Introduction
As shown in Table. 1, this paper aims to evaluate panoramic 3D estimation as a method for generating 3D geometry from the real world and performing lighting analysis. Specifically, this paper makes the following technical contributions:
@@
Table 1: Comparision of Three Approaches for Indoor Lighting Analysis
@@ 3 Methodology
As shown in Fig. 1, three approaches for indoor lighting analysis will be compared in this study. This section presents the details of each approach and simulation settings for generating data.
@@ 3.1 Panoramic HDR Photography
Compared to conventional fisheye images, a single panorama captures the complete content of the space from ${}^{\circ}$ directions. In this study, we focus on a single room and select panoramic HDR photographs at various time points and under different sky conditions from the calibrated panorama dataset [24]. We will compare these calibrated HDR images with panoramic luminance maps generated from computational lighting simulations.
Panoramic photography captures a 3D scene and represents it in an equirectangular representation. In the spherical coordinate system (Fig. 2 (left)), a single point on a unit sphere can be expressed by $\theta$ and $\phi$ (Equation (1)), where $\theta$ represents the Azimuth angle and $\phi$ indicates the Altitude angle. The range of $\theta$ is ($-\pi$, $+\pi$), while the range of $\phi$ is ($-0.5\pi$, $+0.5\pi$).
Meanwhile, given the known $x$, $y$, and $z$ values in the spherical coordinate, the $\theta^{\prime}$ and $\phi^{\prime}$ in equirectangular representation can also be obtained through Equation (2). Fig. 2 (right) shows the mapping process when the 3D point set on the spherical coordinate system is projected onto the equirectangular representation. The transformation between a ${}^{\circ}$ fisheye image and the equirectangular image is invertible. Using the captured indoor panorama, we can modify the azimuth angle to produce fisheye images in different horizontal directions.
@@
$\begin{split}S(\theta,\phi)&=(x,y,z)\\ x&=\cos(\phi)\sin(\theta)\\ y&=\sin(\phi)\\ z&=\cos(\phi)\cos(\theta)\end{split}$ (1)
$\begin{split}\theta^{\prime}&=arctan\frac{x}{z}\\ \phi^{\prime}&=arcsin\frac{y}{\sqrt{x^{2}+y^{2}+z^{2}}}\\ \end{split}$ (2)
@@
To generate 3D room geometry, the captured panorama will be used as an input to estimate indoor wall corners using the LED2Net model [14]. Then, we use the estimated wall corners on the image to segment the panorama and construct the 3D room layout with individual planar surfaces (Fig. 3). Each window, with its unique pattern and geometry, impacts the admission of light from the outdoor to the indoor space. To outline a detailed window pattern, we incorporate the window detection workflow [25]. The modeling approach uses semantic segmentation [26] to identify the window region and line detection [27] to detect the window frames. The window aperture and estimated 3D geometry will be utilized for computational lighting simulation.
@@ 3.3 Detailed Model
The selected space is an existing residential unit (Fig. 4), and the window faces ${}^{\circ}$ South orientation. The detailed scene model was built in Rhino [28] and Grasshopper plugin [29]. Compared to the estimated room geometry with planar surfaces, the detailed model, created through manual inputs, incorporates intricate indoor elements such as door frames, window frames, electrical outlets, smoke detectors, and plumbing, along with wall thickness.
@@ 3.4 Simulation Settings
The selected room is located in Pittsburgh, PA (latitude: ${}^{\circ}$ N, longitude: ${}^{\circ}$ W). The computational lighting simulation is performed in the Radiance-based [8] platform [30], using the local weather file ($_{}^{\_}$$_{}^{\_}$$_{}^{\_}$3.epw) and Radiance parameters (ab: 8, lw: 0.01, samples per pixel: 100). We compare HDR photographs, 3D estimation, and detailed model under two different sky conditions (clear sky and overcast sky). We select time series HDR photographs from July 6th (clear sky) and June 22nd (overcast sky) from the calibrated panorama dataset [24].
Daylight Glare Probability (DGP) is a glare metric that quantifies the percentage of occupants who perceive a scene as glary and quantifies the discomfort glare caused by natural sunlight. Higher DGP values correspond to a greater possibility for glare issues among building occupants [31]. The DGP classifies the result into four glare levels: imperceptible (> 0.35), perceptible (0.35 - 0.40), disturbing (0.40 - 0.45), and intolerable (> 0.45) glare. For DGP analysis, the entire ${}^{\circ}$ view is segmented into sixteen divisions with ${}^{\circ}$ increments (Fig. 4). In the computational simulation for the 3D estimation model and detailed model, the hourly DGP results (from 8:30 to 16:30) from the target dates (March 21st, June 21st, September 21st, and December 21st) are generated from the complete annual glare simulation.
@@ 4.1 Panoramic Luminance Map
The luminance maps from HDR photography, the 3D estimation model, and the detailed model will be compared side by side. Given $R$, $G$, and $B$ values in the captured indoor HDR image, each image is linearly calibrated by $k$ obtained from on-site luminance measurement, and the scene luminance is processed by per-pixel luminance calculation [32] (Equation (3)).
$L=k\cdot(0.2127\cdot R+0.7151\cdot G+0.0722\cdot B)(cd/m^{2})$ (3)
As shown in Fig. 5, we selected the HDR photographs on July 6th under clear sky conditions, with hourly intervals ranging from 07:00 to 14:00. This selection was made considering that direct sunlight cannot be observed from the scene after 14:00. Using the same time intervals, we obtained panoramic luminance maps from computational lighting simulation using the 3D estimation method and the detailed modeling approach, respectively. Compared to the detailed model method, the 3D estimation method brings larger regions of direct sunlight into the scene.
@@ 4.2 Panoramic Error Analysis
The error analysis was conducted on a per-pixel luminance comparison ranging from - 3,000 $cd/m^{2}$ to 3,000 $cd/m^{2}$, where blue indicates underestimation and red represents overestimation, respectively. We focus on clear sky condition on July 6th, 2023 (Fig. 6) and overcast sky condition on June 22nd, 2023 (Fig. 7).
@@
In the clear sky scenario, after the error analysis of the 3D Estimation method and HDR Photography method (Fig. 6 (first column)), it was observed that the 3D estimation method caused higher luminance errors in outdoor scenes, and the ceiling consistently shows overestimation of luminance value, due to the absence of surrounding context in the simulation. For the wall and floor surfaces, minor errors occurred when no direct sunlight came from the outdoor scene. When direct sunlight appears on the indoor surfaces, the planar window surfaces in the 3D estimation model consistently bring extra sunlight region on the wall and floor surfaces.
When comparing the 3D Estimation approach and the Detailed Model approach (Fig. 6 (second column)), the primary luminance differences are located around the window frame area. The 3D estimation method consistently brings larger direct sun regions (over 3,000 cd/$m^{2}$) into the indoor spaces. For other surfaces (like walls, floors, and ceilings), luminance error is small when there is no direct sunlight within the space.
When comparing the Detailed Model approach and HDR Photography approach (Fig. 6 (third column)), similar to digital estimation, the detailed model also exhibits errors in outdoor scenes due to a lack of surrounding context. The luminance value on the the ceiling is also overestimated in computational lighting simulation, due to the absence of outdoor context. The overestimation of the direct sun region on the wall and floor is less in the detailed model method, due to the presence of wall thickness and detailed window frame.
Under overcast sky conditions (Fig. 7), the luminance errors in the indoor space are smaller than in clear sky scenarios (Fig. 6). Without outdoor context in the computational simulation, errors were mainly concentrated on the window region.
@@ 4.3 Fisheye Luminance Map
Figure 8: ${}^{\circ}$ fisheye images in the false color scheme (from 0 to 1,000 cd/$m^{2}$) segmented from HDR photographs, 3D Estimation, and Detailed Models under clear sky conditions (July 6, 2023) and overcast sky conditions (June 22, 2023).
Except for comparing panoramic luminance maps from three methods, Fig. 8 presents the point-in-time results in fisheye perspective under clear sky condition on July 6th, 2023, at 09:56 and overcast sky condition on June 22nd, 2023, at 09:56. The results are displayed in false color ranging from 0 to 1,000 $cd/m^{2}$. For HDR photography, each panoramic image was converted to eight fisheye images with a ${}^{\circ}$ increment. The results of the 3D estimation method and the detailed model method are generated from computational simulation and displayed in ${}^{\circ}$ fisheye perspective side by side.
@@ 4.4 Glare Metric
$\mathrm{DGP}=5.87\times 10^{-5}\mathrm{Ev}+9.8\times 10^{-2}\log\left(1+\sum_{\mathrm{i}}\frac{\mathrm{L}_{\mathrm{s},\mathrm{i}}^{2}\omega_{\mathrm{s},\mathrm{i}}}{\mathrm{Ev}^{1.87}\mathrm{P}_{\mathrm{i}}^{2}}\right)+0.16$ (4)
@@
Computational lighting simulation provides flexibility in calculating the potential glare problems on an annual basis. Given the annual glare simulation, we selected hourly glare results on March 21st, June 21st, September 21st, and December 21st. The target time interval on each day spans hourly from 08:30 to 14:30. As shown in Fig. 4, the view division lists from #1 to #16, which represents the rotating angle from ${}^{\circ}$ to ${}^{\circ}$ in ${}^{\circ}$ gap starting from the view #1.
@@
On March 21st (Fig. 9 (top)), when the window is within the field of view, the DGP shows higher levels around noon time than other time points. Compared to the detailed model, the 3D estimation method results in higher glare levels due to the lack of detailed window structure and wall thickness. On June 21st (Fig. 9 (bottom)) and September 21st (Fig. 10 (top)), only perceptible glare is observed from the detailed model, while higher DGP values in Disturbing glare are generated from the 3D estimation model. On December 21st (Fig. 10 (bottom)), the DGP value exhibits high values when the sun is visible from the window.
@@ 5 Conclusion
After evaluating the luminance maps from captured HDR photographs, the 3D Estimation model, and the Detailed Model, the results demonstrate that the 3D estimation method can accurately reconstruct room geometry. However, indoor scene details are simplified as planar surfaces, and detailed window frames and structures cannot be included in the 3D model.
This research presents panoramic 3D estimation as an approach for generating room geometry to support lighting simulation. By comparing results in panoramic representation and fisheye perspective, it demonstrates that HDR photography, the 3D estimation method, and the detailed model have lower luminance errors on overcast days, whereas more errors occur when direct sunlight enters from the window under clear sky conditions.
6 Limitations and Future Work
This study has potential limitations. First, it only focused on a single room to compare results with different weather conditions and times of the day. Future studies will include other indoor scenes with more complicated spaces, such as kitchens, living rooms, and other functional rooms. Second, only two dates and time-series panoramas were selected from the existing dataset. On-site HDR photography for other dates and time points will be necessary to further compare the luminance error. Third, the existing HDR images were collected in Pittsburgh, USA, and other geographic locations should also be considered.

### v2

# Cheng and Ji

@@ Abstract
Keywords: HDR Photography, 3D Estimation, Panorama, Lighting Simulation.
@@ 1 Introduction
As shown in Table 1, this paper aims to evaluate panoramic 3D estimation as a method for generating 3D geometry from the real world and performing lighting analysis. Specifically, this paper makes the following technical contributions:
@@
Table 1: Comparision of Three Approaches for Indoor Lighting Analysis.
@@ 3 Methodology
As shown in Figure. 1, three approaches for indoor lighting analysis will be compared in this study. This section presents the details of each approach and simulation settings for generating data.
@@ 3.1 Panoramic HDR Photography
Compared to conventional fisheye images, a single panorama captures the complete content of the space from 360 directions. In this study, we focus on a single room and select panoramic HDR photographs at various time points and under different sky conditions from the calibrated panorama dataset [24]. We will compare these calibrated HDR images with panoramic luminance maps generated from computational lighting simulations.
Panoramic photography captures a 3D scene and represents it in an equirectangular representation. In the spherical coordinate system (Figure. 2 (left)), a single point on a unit sphere can be expressed by $\theta$ and $\phi$ (1), where $\theta$ represents the Azimuth angle and $\phi$ indicates the Altitude angle. The range of $\theta$ is ($-\pi$, $+\pi$), while the range of $\phi$ is ($-0.5\pi$, $+0.5\pi$).
Meanwhile, given the known $x$, $y$, and $z$ values in the spherical coordinate, the $\theta^{\prime}$ and $\phi^{\prime}$ in equirectangular representation can also be obtained through (2). Figure. 2 (right) shows the mapping process when the 3D point set on the spherical coordinate system is projected onto the equirectangular representation. The transformation between a 180 fisheye image and the equirectangular image is invertible. Using the captured indoor panorama, we can modify the azimuth angle to produce fisheye images in different horizontal directions.
@@
$\begin{split}S(\theta,\phi)&=(x,y,z);\\ x&=\cos(\phi)\sin(\theta);\\ y&=\sin(\phi);\\ z&=\cos(\phi)\cos(\theta).\end{split}$ (1)
$\begin{split}\theta^{\prime}&=arctan\frac{x}{z};\\ \phi^{\prime}&=arcsin\frac{y}{\sqrt{x^{2}+y^{2}+z^{2}}}.\\ \end{split}$ (2)
@@
To generate 3D room geometry, the captured panorama will be used as an input to estimate indoor wall corners using the LED2Net model [14]. Then, we use the estimated wall corners on the image to segment the panorama and construct the 3D room layout with individual planar surfaces (Figure. 3). Each window, with its unique pattern and geometry, impacts the admission of light from the outdoor to the indoor space. To outline a detailed window pattern, we incorporate the window detection workflow [25]. The modeling approach uses semantic segmentation [26] to identify the window region and line detection [27] to detect the window frames. The window aperture and estimated 3D geometry will be utilized for computational lighting simulation.
@@ 3.3 Detailed Model
The selected space is an existing residential unit (Figure. 4), and the window faces 167 South orientation. The detailed scene model was built in Rhino [28] and Grasshopper plugin [29]. Compared to the estimated room geometry with planar surfaces, the detailed model, created through manual inputs, incorporates intricate indoor elements such as door frames, window frames, electrical outlets, smoke detectors, and plumbing, along with wall thickness.
@@ 3.4 Simulation Settings
The selected room is located in Pittsburgh, PA (latitude: 40.4406 N, longitude: 79.9959 W). The computational lighting simulation is performed in the Radiance-based [8] platform [30], using the local weather file (USA_PA_Pittsburgh.Intl.AP.725200_TMY3.epw) and Radiance parameters (ab: 8, lw: 0.01, samples per pixel: 100). We compare HDR photographs, 3D estimation, and detailed model under two different sky conditions (clear sky and overcast sky). We select time series HDR photographs from July 6th (clear sky) and June 22nd (overcast sky) from the calibrated panorama dataset [24].
Daylight Glare Probability (DGP) is a glare metric that quantifies the percentage of occupants who perceive a scene as glary and quantifies the discomfort glare caused by natural sunlight. Higher DGP values correspond to a greater possibility for glare issues among building occupants [31]. The DGP classifies the result into four glare levels: imperceptible (> 0.35), perceptible (0.35 - 0.40), disturbing (0.40 - 0.45), and intolerable (> 0.45) glare. For DGP analysis, the entire 360 view is segmented into sixteen divisions with 22.5 increments (Figure. 4). In the computational simulation for the 3D estimation model and detailed model, the hourly DGP results (from 8:30 to 16:30) from the target dates (March 21st, June 21st, September 21st, and December 21st) are generated from the complete annual glare simulation.
@@ 4.1 Panoramic Luminance Map
The luminance maps from HDR photography, the 3D estimation model, and the detailed model will be compared side by side. Given $R$, $G$, and $B$ values in the captured indoor HDR image, each image is linearly calibrated by $k$ obtained from on-site luminance measurement, and the scene luminance is processed by per-pixel luminance calculation [32] (3).
$L=k\cdot(0.2127\cdot R+0.7151\cdot G+0.0722\cdot B)(cd/m^{2}).$ (3)
As shown in Figure. 5, we selected the HDR photographs on July 6th under clear sky conditions, with hourly intervals ranging from 07:00 to 14:00. This selection was made considering that direct sunlight cannot be observed from the scene after 14:00. Using the same time intervals, we obtained panoramic luminance maps from computational lighting simulation using the 3D estimation method and the detailed modeling approach, respectively. Compared to the detailed model method, the 3D estimation method brings larger regions of direct sunlight into the scene.
@@ 4.2 Panoramic Error Analysis
The error analysis was conducted on a per-pixel luminance comparison ranging from - 3,000 $cd/m^{2}$ to 3,000 $cd/m^{2}$, where blue indicates underestimation and red represents overestimation, respectively. We focus on clear sky condition on July 6th, 2023 (Figure. 6) and overcast sky condition on June 22nd, 2023 (Figure. 7).
@@
In the clear sky scenario, after the error analysis of the 3D Estimation method and HDR Photography method (Figure. 6 (first column)), it was observed that the 3D estimation method caused higher luminance errors in outdoor scenes, and the ceiling consistently shows overestimation of luminance value, due to the absence of surrounding context in the simulation. For the wall and floor surfaces, minor errors occurred when no direct sunlight came from the outdoor scene. When direct sunlight appears on the indoor surfaces, the planar window surfaces in the 3D estimation model consistently bring extra sunlight region on the wall and floor surfaces.
When comparing the 3D Estimation approach and the Detailed Model approach (Figure. 6 (second column)), the primary luminance differences are located around the window frame area. The 3D estimation method consistently brings larger direct sun regions (over 3,000 cd/$m^{2}$) into the indoor spaces. For other surfaces (like walls, floors, and ceilings), luminance error is small when there is no direct sunlight within the space.
When comparing the Detailed Model approach and HDR Photography approach (Figure. 6 (third column)), similar to digital estimation, the detailed model also exhibits errors in outdoor scenes due to a lack of surrounding context. The luminance value on the the ceiling is also overestimated in computational lighting simulation, due to the absence of outdoor context. The overestimation of the direct sun region on the wall and floor is less in the detailed model method, due to the presence of wall thickness and detailed window frame.
Under overcast sky conditions (Figure. 7), the luminance errors in the indoor space are smaller than in clear sky scenarios (Figure. 6). Without outdoor context in the computational simulation, errors were mainly concentrated on the window region.
@@ 4.3 Fisheye Luminance Map
Figure 8: 180 fisheye images in the false color scheme (from 0 to 1,000 cd/$m^{2}$) segmented from HDR photographs, 3D Estimation, and Detailed Models under clear sky conditions (July 6, 2023) and overcast sky conditions (June 22, 2023).
Except for comparing panoramic luminance maps from three methods, Figure. 8 presents the point-in-time results in fisheye perspective under clear sky condition on July 6th, 2023, at 09:56 and overcast sky condition on June 22nd, 2023, at 09:56. The results are displayed in false color ranging from 0 to 1,000 $cd/m^{2}$. For HDR photography, each panoramic image was converted to eight fisheye images with a 45 increment. The results of the 3D estimation method and the detailed model method are generated from computational simulation and displayed in 180 fisheye perspective side by side.
@@ 4.4 Glare Metric
$\mathrm{DGP}=5.87\times 10^{-5}\mathrm{Ev}+9.8\times 10^{-2}\log\left(1+\sum_{\mathrm{i}}\frac{\mathrm{L}_{\mathrm{s},\mathrm{i}}^{2}\omega_{\mathrm{s},\mathrm{i}}}{\mathrm{Ev}^{1.87}\mathrm{P}_{\mathrm{i}}^{2}}\right)+0.16.$ (4)
@@
Computational lighting simulation provides flexibility in calculating the potential glare problems on an annual basis. Given the annual glare simulation, we selected hourly glare results on March 21st, June 21st, September 21st, and December 21st. The target time interval on each day spans hourly from 08:30 to 14:30. As shown in Figure. 4, the view division lists from #1 to #16, which represents the rotating angle from 0 to 360 in 22.5 gap starting from the view #1.
@@
On March 21st (Figure. 9 (top)), when the window is within the field of view, the DGP shows higher levels around noon time than other time points. Compared to the detailed model, the 3D estimation method results in higher glare levels due to the lack of detailed window structure and wall thickness. On June 21st (Figure. 9 (bottom)) and September 21st (Figure. 10 (top)), only perceptible glare is observed from the detailed model, while higher DGP values in Disturbing glare are generated from the 3D estimation model. On December 21st (Figure. 10 (bottom)), the DGP value exhibits high values when the sun is visible from the window.
@@ 5 Conclusion
After evaluating the luminance maps from captured HDR photographs, the 3D Estimation model, and the Detailed Model, the results demonstrate that the 3D estimation method can accurately reconstruct room geometry. However, indoor scene details are simplified as planar surfaces, and detailed window frames and structures cannot be included in the 3D model. This research presents panoramic 3D estimation as an approach for generating room geometry to support lighting simulation. By comparing results in panoramic representation and fisheye perspective, it demonstrates that HDR photography, the 3D estimation method, and the detailed model have lower luminance errors on overcast days, whereas more errors occur when direct sunlight enters from the window under clear sky conditions.
6 Limitation and Future Work
This study has several limitations. First, it only focused on a single room to compare results with different weather conditions and times of the day. Future studies will include other indoor scenes with more complicated spaces, such as kitchens, living rooms, and other functional rooms. Second, only two dates and time-series panoramas were selected from the existing dataset. On-site HDR photography for other dates and time points will be necessary to further compare the luminance error. Third, the existing HDR images were collected in Pittsburgh, USA, and other geographic locations should also be considered.
Author Biographies
ZINING CHENG is an Architecture Designer at Perkins&Will, USA. Her research interests are modeling and simulation with applications in optimizing the indoor built environment for energy efficiency and human comfort. Her email address is .
GUANZHOU JI is a PhD candidate at Carnegie Mellon University, USA. His research interests are indoor photometry, computational photography, and scene editing. His email address is .